A turkish composite bow can release an arrow with a speed of 100 m/s.

(a) if a turk were to fire the bow at 10 degrees above the horizontal in order to hit someone lying 5 metres below him, if air resistance is negligible, what would be:

(i) the time taken?

(ii) the horizontal distance x?

(b) repeat the above calculations assuming vertical air resistance is negligible, but horizontally it causes a deceleration of 0.5 m/s2.


1
Expert's answer
2021-11-25T16:23:38-0500

a)

"t_1=\frac{v\sin\alpha}g=1.77~s,"

"s_1=\frac{v^2\sin2\alpha}{2g}=174.5~m,"

"t_2=\frac{\sqrt{v^2\sin^2\alpha+2gh}}g=2.04~s,"

"s_2=\frac{v\cos\alpha}g\sqrt{v^2\sin^2\alpha+2gh}=200.9~m,"

"t=t_1+t_2=3.81~s,"

"s=s_1+s_2=375.4~m."

b)

"t_1'=t_1=1.77~s,"

"s_1'=s_1-\frac{at_1^2}2=173.7~m,"

"t_2'=t_2=2.04~m,"

"s_2'=s_2-\frac{at_2^2}2=199.9~m,"

"t'=t=3.81~s,"

"s'=s_1'+s_2'=373.6~m."